OpenStudy (anonymous):

Derivative of x^3*e^(2x) I get different results when I check my answer to Wolfram Alpha and derivativecalculator.net Please help, it should be easy.

ganeshie8 (ganeshie8):

use chain rule

ganeshie8 (ganeshie8):

and product rule

ganeshie8 (ganeshie8):

\(\large \mathbb [x^3*e^{2x} ]' = x^3 [e^{2x}]' + [x^3]' e^{2x}\)

OpenStudy (anonymous):

What? f´(x) = \[\frac{ d }{ dx }\left( x^3 \right) * e^(2x) + x^3* \frac{ d }{ dx }\left( e ^{2x}\right)\]

ganeshie8 (ganeshie8):

looks good^

OpenStudy (anonymous):

\[suppose~ y=x^3e^(2x) \] \[if ~y=uv,y'=uv'+u'v\]

OpenStudy (anonymous):

Then I get: \[3x^2 * e ^{2x} + x^3* 2e ^{2x}\] which seems to be wrong...or?

ganeshie8 (ganeshie8):

It is correct !

ganeshie8 (ganeshie8):

did wolfram give a different answer ?

OpenStudy (anonymous):

but not according to wolfram alpha...though. Maybe I did it wrong...

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

wolfram just factored the answer

OpenStudy (anonymous):

ot may be\[x^2e ^{2x}\left( 2x+3 \right)\]

ganeshie8 (ganeshie8):

factor out the GCF in your answer

OpenStudy (anonymous):

Aaaah, now I get it...if was factored ,yes @ganeshie8. Thank you all. I don't seem to know what I'm doing haha.

ganeshie8 (ganeshie8):

:) you're doing it correctly.... and yes its always good to double/triple confirm wid wolfram !